The value of ΔG°\' for the conversion of dihydroxyacetone phosphate to glyceraldehyde-3-phosphate (GAP) is 7.53 kJ/mol. If the concentration of dihydroxyacetone phosphate at equilibrium is 2.85 mM, what is the concentration of glyceraldehyde-3-phosphate? Assume a temperature of 25.0 °C. The constant R = 8.3145 J/(mol·K)Map do Sapling Learning macmilan learning The value of AG° for the conversion of dihydroxyacetone phosphate to glyceraldehyde
Answer #1

Dear student

Use the following equation and solve for K:

e^∆G/-RT=K

Where

∆G=7.53kJ/mol=7.53 x 10^3 J/mol
R=8.3145 J/(mol·K)
T=273K + Cº=273K+25=298K

e^[7.53 x 10^3 J/mol/-8.3145 J/(mol·K)*298K]=K

K=4.788 x 10^-2

The reaction is as followed:

dihydroxyacetone phosphate -------------------> glyceraldehyde-3-phosphate


K=products/reactants

K=[glyceraldehyde-3-phosphate]/[dihydroxyacetone phosphate]

K=x/2.35mM

K=4.788 x 10^-2=x/2.35mM

Solve for x:

(4.788 x 10^-2)*(2.35mM)=x

x=0.112mM

glyceraldehyde-3-phosphate=0.112mM at equilibrium

Regards

Team Keep Learning

Please do upvote if you like my answer